# HeadCraft: Modeling High-Detail Shape Variations for Animated 3DMMs

### v1

@@ 4.2 Results
Unconditional sampling. In Fig. 3, we compare the difference in details and diversity of the unconditional samples produced by our method to the ones produced by NPHM [26] and ROME [32] methods. Additionally, we compare it to the PCA baseline, whereas PCA linear basis is fitted to our UV displacement maps, and provide the numbers for random FLAME samples without added displacements as a reference. For Ours, PCA baseline and ROME, a FLAME with random shape, expression and jaw parameters are sampled from a Gaussian for every head mesh, in accordance with the statistics precalculated over the NeRSemble dataset [34]. While NPHM, PCA baseline, and Ours have been fitted to exactly the same training dataset, for ROME, the authors’ provided checkpoint has been used. For ROME, we sample the FLAME displacements from a linear model provided by the authors of ROME as the sampling strategy proposed by the ROME authors. Visually, we observe both higher diversity and better representation of hair details than for all baselines. The details of the facial region are generally the sharpest for ROME, PCA baseline, and Ours, due to the use of the FLAME template.
In Table 1, we also quantify the level of detail and variety of the generated meshes w.r.t. the full head scans from the FaceVerse dataset [55] that has not been used for training. The comparison is performed in two ways.
FID ↓ KID ↓ IS ↑ Rel. IS ↑ MMD ↓ JSD ↓ COV ↑
Ours 72.37 0.071 1.67 88.25% 6.45 21.41 47.12%
PCA 102.96 0.125 1.46 77.13% 9.96 20.68 22.12%
NPHM [26] 139.82 0.170 1.57 82.95% 7.80 19.06 46.15%
ROME [32] 169.65 0.204 1.63 86.81% 10.02 23.19 32.69%
FLAME [36] 198.85 0.262 1.13 59.86% 12.95 23.89 5.77%
Table 1: The comparison of quality and diversity of random samples generated by each of the methods. FID and KID measure the similarity of the generated mesh renderings vs. the renderings of the ground truth meshes in FaceVerse dataset, while IS measures the realism of the generated renderings. 3D metrics MMD, JSD, COV assess the similarity of point clouds sampled from generated and ground truth meshes. For this and all other tables, MMD has been multiplied by $10^{3}$ and JSD is multiplied by $10^{2}$.
Firstly, to evaluate the visual plausibility of the generated geometry, we render 2195 ground truth meshes from FaceVerse and the same number of meshes generated by each method with highly metallic material from eight distinct viewpoints, uniformly sampled along the circular trajectory in the horizontal plane. The FID [28] and KID [9] perceptual metrics are calculated for all generated and ground truth renderings from a given viewpoint and then macro-averaged over eight viewpoints. Inception Score (IS[47] evaluated the realism of the generated renderings. Because of that, we also report the relative IS as a percentage of the reference IS calculated for the ground truth renderings.
Secondly, we compare the distributions of point clouds sampled from the generated and ground truth meshes. To do that, we sample 10K points from each of the 2195 generated and the same number of ground truth meshes and calculate several 3D similarity metrics. Jensen-Shannon Divergence (JSD) is evaluated by comparing the distributions of generated and ground truth points, splat into a voxel grid (in our case, of $512^{3}$ voxels). Minimum Matching Distance (MMD) is a measure of 3D object realism that, for each ground truth sample, involves evaluating the distance to the most similar sample in the generated set. Similarly, Coverage (COV) indicates the percentage of the ground truth samples, for which the nearest neighbor among all ground truth and generated samples falls into the generated set. The detailed mathematical description of each of the 3D metrics we report can be found in [58, 21]. In addition, we demonstrate how much the generated samples deviate from the NPHM training set in Fig. 5. The nearest neighbor scan is found by comparing the generated UV map to the UV maps comprised of registered ground truth displacements for all training scans by L2 distance over the scalp region.
@@
Generated Nearest Generated Nearest
@@
Figure 5: Randomly generated samples from HeadCraft and the corresponding nearest neighbors in the NPHM dataset among the scans used for training.
    Standard UV [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
      Custom UV [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 6: Ablation over the choice of the UV layout. Our method utilizes custom UV layout that allows us to model more consistent geometry by mitigating seam artifacts, as seen from the back view here. Layouts are shown in Fig. 4.
Ablating over the choice of the generative model architecture. We compare StyleGAN to other state-of-the-art generative model architectures, namely of VAE [33] and VQ-VAE [53] family, with ResNet-18 encoder and decoder. For VQ-VAE, the sampling from the latent space is implemented via training PixelCNN autoregressive model [52]. The results are presented in Table 2 and in Fig. 7.
FID ↓ KID ↓ IS ↑ Rel. IS ↑ MMD ↓ JSD ↓ COV ↑
Ours 72.37 0.07 1.67 88.25% 6.45 21.41 47.12%
VAE [33] 108.91 0.13 1.66 87.79% 6.79 21.65 42.31%
VQ-VAE [53] 125.65 0.15 1.62 85.94% 7.96 21.88 38.46%
PCA 102.96 0.13 1.46 77.13% 9.96 20.68 22.12%
@@ 4.3 Applications
Fitting the latent code to a depth map. Our model can act as a prior for completing the partial observations, e.g. when they come from a depth sensor. To evaluate the performance of the model in that scenario, we demonstrate the completion capabilities of the model over a number of scans from NPHM corresponding to the subjects unseen during training. For each of these scans, we project their depth onto random viewpoints in the frontal hemisphere and project it back to 3D to construct partial point clouds. To obtain a partial UV map to be completed, we run our registration procedure with a few modifications to fit a part of the scan. Namely, we only fit the points within the convex hull of the partial point cloud, apply stronger edge length regularization weight, and constrain the points at the border of the allowed region from moving. The final mask of observed UV texels is refined by only selecting those points that turn out to be close to the partial point cloud. Finally, a latent code of HeadCraft explaining the partial UV map is found via StyleGAN inversion techniques. More technical details of the partial registration and inversion are provided in the Supplementary [3]. The fitting quality can be evaluated by the visual comparison in Fig. 8.
VAE VQ-VAE
Figure 7: Ablation over the generative model design. VAE and VQ-VAE both follow the ResNet-18 encoder and decoder architecture, while our method is based on StyleGAN2. The results from VAE and VQ-VAE match the diversity of the training data but not the level of detail and handle the UV seam worse.
@@
The capabilities of fitting the model to the full scan, e.g. created from Structure-from-Motion (SfM), are demonstrated as a part of the semantic editing experiments in Fig. 11 (top rows, the result of the latent fitting to each of the scans, $\lambda=\{0,1\}$).
Animation. The decomposition of the parametric model and the displacements allows us to animate the complete head model. In our experiments, we take real multi-view video sequences with talking people from the NeRSemble dataset [34] and obtain shape, expression, jaw, and head pose parameters for each time frame of the speech by running a FLAME tracker for each sequence. For each of the sample shapes, estimated from the sequences, we reenact the corresponding FLAME with estimated expression parameters, subdivide the template and query a randomly pre-sampled UV displacement map from HeadCraft. Since the template is also deforming over time, we rotate the displacements according to the changing surface normals of the template. More results are demonstrated in Fig. 9 and in the Supplementary Video.
Figure 9: Demonstration of the animation capabilities of the model. Each of the sequences is created by taking FLAME shape parameters, expression, jaw, and head pose parameters from a sequence from NeRSemble dataset [34], subdividing the template, and applying randomly generated displacements from HeadCraft.
(a) Stage 1 (b) Stage 2 (c) Stage 1, (d) Ours
only only $\boldsymbol{\lambda}=\boldsymbol{\lambda}_{\textrm{Stage 2}}$
Figure 10: Ablation over the one-stage vs. two-stage registration. Leaving only the vector displacements stage (a) yields too smooth geometry, and learning them only along the normals (b) introduces unnecessary spikes – just like running the first stage with weaker regularization (c).
4.4 Analysis
@@
Figure 11: Semantic editing. Interpolating the latent representations of HeadCraft allows us to smoothly change the person’s appearance from one to another. To do that, we fit the latent codes for two real scans in the NPHM dataset (brown, top rows), unseen during training, and blend them together with a $\lambda$ weight. Likewise, we can transfer hair geometry from one person to another.
Hair transfer from one scan to another. Access to the shared UV space allows us to modify the geometry semantically. In Fig. 11, the transfer of the scalp region from one ground truth NPHM scan, unseen during training, to the other is shown. The transfer is performed via fitting the latent representation of HeadCraft to the driver scan (the source of displacements) and feeding it to the model. The extracted displacements are later applied to the source scan.
@@ 5 Discussion
In this work, a generative model for 3D human heads is presented. We demonstrate the efficacy of the hybrid approach involving an underlying animatable parametric model and a neural vertex displacement modeller. Most importantly, our method allows to model high-quality shape variations while maintaining the realistic animation capability, and the inversion framework allows us to find a suitable latent representation to either represent a full head scan or a part of it that could come from e.g. the depth sensor. A direction of the possible future work could be focused on incorporating an appearance model for color and material-based relighting and a physical model of hair movement, based on, for instance, hair strands, to support more realistic animation.
Acknowledgments.
We gratefully acknowledge the support of this research by a TUM-IAS Hans Fischer Senior Fellowship, the ERC Starting Grant Scan2CAD (804724) and the Horizon Europe vera.ai project (101070093). We also thank Yawar Siddiqui and Alexey Artemov for helpful advice, Tobias Kirschstein for his assistance with the NeRSembles dataset and visualization, Taras Khakhulin for his help with the ROME baseline, Peter Kocsis, Antonio Allegro, Jiapeng Tang for early peer review, Silvia Sellán for the Blender visualization course materials that we used, and Angela Dai for the video voiceover.
@@ A.1 Displacement registration procedure
During the vector displacement stage, only the scalp region (defined by the semantic mask shipped with FLAME [36]) is optimized. During the normal displacement stage, we also unfreeze the facial region but keep the neck, eyeballs, ears and inner mouth region frozen (the latter is annotated manually in Blender [18] package and is frozen because of its absence in the ground truth scans, as it is placed fully interior). Each stage takes around 3 min for 1K steps on a single NVIDIA RTX 2080 Ti GPU. We used the PyTorch3D [44] functions for implementation of all the loss terms.
In Fig. 12, we show how the seam was annotated for the custom UV layout. The Blender [18] 4.0 package was used for annotation.
Standard UV Custom UV
[Uncaptioned image] [Uncaptioned image]
Figure 12: Demonstration of the custom UV seam annotation. Standard UV corresponds to the UV space shipped with the FLAME model. Custom UV stands for the handcrafted layout employed in our pipeline. This layout simplifies learning sophisticated shape variations such as long hair in a consistent way.
@@ A.2 Generative model
The model was trained on four NVIDIA RTX 2080 Ti GPUs. For training of the StyleGAN, we replaced the facial part of the registered UV maps of subjects in the NPHM dataset with the corresponding facial part from the neutral expression scan of the same person. This has been done to better support various expressions We also found it beneficial to disable the StyleGAN noise, typically injected into the generator, for the face part of the UV map, to smooth out the generated facial displacements relative to the scalp displacements that normally require higher level of detail. For the scalp region, the StyleGAN noise is constant, initialized separately for each generator layer before training.
Post-processing. The non-smoothness of the seam is mitigated in two ways, which are described in more detail here. The first step is blending the displacements. To do that, we analyze the UV coordinates of the triangles in the custom layout $\mathcal{U}_{\mathcal{F}}$ and find the triangles, for which at least one of its vertices lies on the seam. These are the triangles, for which some of the UV coordinates belong to the left part of the UV map (scalp) and some belong to the right part (face). This also provides us with the texel (i.e. UV map pixel) correspondence between the seam vertices. To softly blend the vertices inside the scalp region (left part of the UV map) near the seam, we first copy the displacements from the seam texels on the right to the respective texels on the left. Then, we compute the coordinates of the respective nearest seam texel from all texels inside the scalp region by applying distance transform algorithm [48]. In the vicinity of 10 texels from the seam (in any direction inside the region), we softly blend the displacements between the original and the seam ones:
$U[i,j]$ $:=\frac{10-\textrm{dist\_to\_seam}[i,j]}{10}U[\textrm{nn}_{i}[i,j],\textrm{nn}_{j}[i,j]]$
$+\frac{\textrm{dist\_to\_seam}[i,j]}{10}U[i,j],$
[…]
The second step is aimed to fix the sampling errors that can occur when sampling a generated UV map. To do that, we again calculate the distance transform, this time for the empty space in the UV map (regions corresponding to no useful information). This way, the nearest to the seam texel is found for each texel of the empty space, and we copy the displacement from the nearest seam texel into this texel. Effectively, this replaces the empty space with the Voronoi diagram [24] of proximity to the seam, with displacements from the seam occupying each region of the diagram. As a result, when the displacements are being sampled from the UV map with the empty space filling procedure applied in advance, no UV coordinates fall into the space where no displacements are set. In case of hitting the space that was previously empty, the displacement attains a value close to the nearest displacement at the seam.
Appendix B Results: Technical Details
@@
Unconditional sampling. In Fig. 13, we provide more unconditional samples from our model from different viewpoints. All the samples have been produced by sampling $z\in\mathcal{N}(0,\mathbb{I})$ with a truncation trick [11] with the power $\psi=0.7$. For the evaluation in the Table 1 in the main text, the implementation of the metrics MMD, JSD, COV from PointFlow [58] was used. Since FaceVerse contains samples grouped by subjects, the nearest neighbor of a ground truth scan is typically a scan of the same subject with a different expression. Because of that, we only select one ground truth sample per subject (with the same neutral expression for all subjects) to calculate COV. All FaceVerse scans are used to calculate MMD and JSD. As a distance measure between individual point clouds, aggregated over multiple observations in MMD and COV, we use Chamfer Distance (CD).
@@
Behavior of the registration procedure. In Fig. 14, 15, 16, 17, we show how the mesh deforms as a result of the vector displacements optimization and normal displacements optimization.

### v2

@@ 4.4 Results
Unconditional sampling. In Fig. 3, we compare the difference in details and diversity of the unconditional samples produced by our method to the ones produced by NPHM [24] and ROME [32] methods. For Ours, PCA baseline and ROME, a FLAME with random shape, expression and jaw parameters are sampled from normal distribution for every head mesh, in accordance with the statistics precalculated over the NeRSemble dataset [34]. For ROME, we sample the FLAME displacements from a linear model provided by the authors of ROME as the sampling strategy proposed by the ROME authors. Visually, we observe both higher diversity and better representation of details than for all baselines. The details of the facial region are generally the sharpest for ROME, PCA baseline, and Ours, due to the use of the FLAME template.
In Table 1, we also quantify the level of detail and variety of the generated meshes w.r.t. the full head scans from the FaceVerse dataset [56] that has not been used for training. In addition, we demonstrate how much the generated samples deviate from the NPHM training set in Fig. 4. Nearest is found by comparing the generated displacement map to the maps of registered ground truth displacements for all training scans by L2 distance over the scalp.
@@
Generated Nearest Generated Nearest Generated Nearest
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 4: Randomly generated samples from HeadCraft and the corresponding nearest neighbors in the NPHM dataset among the scans used for training. $L_{2}$ distance over the scalp part of the displacement maps was used. Displacements were added to a random FLAME template for all samples.
Ablating over the choice of the generative model architecture. We compare StyleGAN to other state-of-the-art generative model architectures, namely of VAE [33] and VQ-VAE [54] family, with ResNet-18 encoder and decoder. For VQ-VAE, the sampling from the latent space is implemented via training PixelCNN autoregressive model [53]. The results are presented in Table 2 and can be visually assessed in the Supplementary.
FID ↓ KID ↓ MMD ↓ JSD ↓ COV ↑
Ours 68.00 0.065 6.51 21.33 53.85%
SG $\rightarrow$ VAE 112.03 0.130 6.78 21.67 47.12%
SG $\rightarrow$ VQ-VAE 124.17 0.151 7.15 21.93 43.27%
PCA 126.31 0.165 10.47 20.16 23.08%
   (a) Stage 1    (b) Stage 2    (c) Stage 1,      (d) Ours
only only $\boldsymbol{\lambda}=\boldsymbol{\lambda}_{\textrm{Stage 2}}$
@@
Figure 5: Ablation over the one-stage vs. two-stage registration. Regressing only vector displacements (a) yields too smooth geometry, and learning them only along the normals (b) introduces spikes – just like running the first stage with smaller $\boldsymbol{\lambda}$ (c).
@@
Figure 6: Demonstration of the animation capabilities of the model. Each of the sequences is created by adding the randomly generated displacements from HeadCraft to the FLAME template with varying extreme expression parameters.
@@ 4.5 Applications
Fitting the latent code to a depth map. Our model can act as a prior for completing the partial observations, e.g. when they come from a depth sensor. To evaluate the performance of the model in that scenario, we demonstrate the completion capabilities of the model over a number of scans from NPHM corresponding to the subjects unseen during training. For each of these scans, we project their depth onto random viewpoints in the frontal hemisphere and project it back to 3D to construct partial point clouds. To obtain a partial UV map to be completed, we run our registration procedure with a few modifications to fit a part of the scan. Namely, we only fit the points within the convex hull of the partial point cloud, apply stronger edge length regularization weight, and constrain the points at the border of the allowed region from moving. The final mask of observed UV texels is refined by only selecting those points that turn out to be close to the partial point cloud. Finally, a latent code of HeadCraft explaining the partial UV map is found via StyleGAN inversion techniques. More technical details of the partial registration and inversion are provided in the Supplementary. The fitting quality can be evaluated by the visual comparison in Fig. 7.
@@
Figure 8: Semantic editing. Interpolating the latent representations of HeadCraft (left) allows us to smoothly change the person’s appearance from one to another. To do that, we fit the latent codes for two real scans in the NPHM dataset (brown, top rows), unseen during training, and blend them together with a $\lambda$ weight. Likewise, we can transfer surface variations from one person to another (right). The source person is marked with $\star$ on the left and the driving person is in the bottom row on the right.
The capabilities of fitting the model to the full scan, e.g. created from Structure-from-Motion (SfM), are demonstrated as a part of the semantic editing experiments in Fig. 8 (top rows, the result of the latent fitting, $\lambda=\{0,1\}$).
Animation. The decomposition of the parametric model and the displacements allows us to animate the complete head model. In our experiments, we take real multi-view video sequences with talking people from the NeRSemble dataset [34] and obtain shape, expression, jaw, and head pose parameters for each time frame of the speech by running a FLAME tracker for each sequence. For each of the sample shapes, estimated from the sequences, we reenact the corresponding FLAME with estimated expression parameters, subdivide the template and query a randomly pre-sampled UV displacement map from HeadCraft. Since the template is also deforming over time, we rotate the displacements according to the changing surface normals of the template. The reenactment results on NeRSemble are available in the Supplementary Video. In Fig. 6, for higher clarity, we demonstrate rigging with randomly sampled FLAME shapes and a small number of extreme expressions generated artificially by randomly setting a subset of the first ten expression components of FLAME to ±$\pm 2$.
@@
Semantic transfer from one scan to another. Access to the shared UV space allows us to modify the geometry semantically. In Fig. 8, the transfer of the scalp region from one ground truth NPHM scan, unseen during training, to the other is shown. The transfer is performed via fitting the latent representation of HeadCraft to the driver scan (the source of displacements) and feeding it to the model. The extracted displacements are later applied to the source scan.
@@ 5 Discussion
In this work, a generative model for 3D human heads is presented. We demonstrate the efficacy of the hybrid approach involving an underlying animatable parametric model and a neural vertex displacement modeller. Most importantly, our method allows to model high-quality shape variations while maintaining the realistic animation capability, and the inversion framework allows us to find a suitable latent representation to either represent a full head scan or a part of it that could come from e.g. the depth sensor. A direction of the possible future work could be focused on incorporating an appearance model for color and material-based relighting and a physical model of hair movement, based on, for instance, hair strands, to support more realistic animation. The code and the dataset of displacement registrations will be released to the public.
Acknowledgments. We gratefully acknowledge the support of this research by a TUM-IAS Hans Fischer Senior Fellowship, the ERC Starting Grant Scan2CAD (804724) and the Horizon Europe vera.ai project (101070093). We also thank Yawar Siddiqui, Alexey Artemov, Justus Thies for helpful advice, Tobias Kirschstein for his assistance with the NeRSemble, Taras Khakhulin for his help with the ROME baseline, Angela Dai for the video voiceover.
@@ A.1 Displacement registration procedure
During the vector displacement stage, only the scalp region (defined by the semantic mask shipped with FLAME [36]) is optimized. During the normal displacement stage, we also unfreeze the facial region but keep the neck, eyeballs, ears, and inner mouth region frozen (the latter is annotated manually in Blender [15] package and is frozen because of its absence in the ground truth scans, as it is placed fully interior). Each stage takes around 3 min for 1K steps on a single NVIDIA RTX 2080 Ti GPU. We used the PyTorch3D [48] functions for implementation of all the loss terms.
In Fig. 9, we show how the seam was annotated for the custom UV layout.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 9: Comparison of the FLAME layouts. The standard, commonly used unwrapping for FLAME (left) features a seam corresponding to the vertical line in the back of the head and pays more attention to the facial region than to the scalp. In the hand-crafted custom layout that we employ (right), a different seam around the face border is selected, thus making the regions of face and scalp separated and of similar size to avoid unnecessary breaks and expand the region of interest. Lower neck is not modeled by our method and hence made partially unseen in UV.
Blender [15] 4.0 package was used to annotate the seam for the custom UV layout. The seam is manually annotated to be symmetric w.r.t. the reflection symmetry of FLAME. Firstly, the layout was constructed via ”Unwrap“ Blender tool and adjusted manually via local translation and scaling tools to better fit the available space. After that, the UVs for vertices in the left part of the head were assigned with the mirrored UVs of the vertices in the right part to account for symmetry. This is repeated for the face and scalp separately. Finally, the face, scalp, and eyeballs parts were scaled and translated to fit the unit square. Regions, for which displacements are not modeled, are either given smaller space in the layout (eyeballs) or moved outside of the unit square (unmodeled parts of the neck).
@@
Standard UV [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Custom UV [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 10: Ablation over the choice of the UV layout. Our method utilizes custom UV layout that allows us to model more consistent geometry by mitigating seam artifacts.
@@ A.2 Generative model
The model was trained on four NVIDIA RTX 2080 Ti GPUs. For training of the StyleGAN, the ground truth UV displacement maps were transformed from $256\times 256\times 3$ into $256\times 256\times 6$ by first bilinearly upscaling the displacement maps over the width dimension (resulting in the map of shape $256\times 512\times 3$) and then splitting it into two three-channel maps by the width dimension. This effectively means that StyleGAN predicts the face and scalp as two separate three-channel images, which increases the spatial area in the output dedicated to each region. Additionally, we replaced the facial part of the registered UV maps of subjects in the NPHM dataset with the corresponding facial part from the neutral expression scan of the same person. This has been done to better support various expressions at the inference time. We also found it beneficial to disable the StyleGAN noise at the inference time, typically injected into the generator, for the face part of the UV map to smooth out the generated facial displacements relative to the scalp displacements that normally require a higher level of detail. For the scalp region, the StyleGAN noise is constant, initialized separately for each generator layer before training. This noise schedule separation is performed via two passes through StyleGAN – one with zero noise and another with constant noise – and combining the corresponding results.
Post-processing. To ensure a smooth transition from the face to the scalp part around the seam, we first create a mesh by querying the UV map as is, then identify the seam vertices and apply Laplacian smoothing [55] to the mesh in the $K$-vertex vicinity of the seam (vicinity obtained via BFS-style expansion of the seed vertices defining the seam in a graph defined by the mesh edges). In our experiments, $K=10$ (corresponding to the subdivided template with roughly 100K vertices), and Laplacian smoothing is repeated 10 times, only affecting the seam vicinity area. Since the face displacements are modeled with lower smoothness due to the disabled StyleGAN noise in that area, the post-processing technique results in a smooth transition from the relatively smooth face displacements to sharper details in the scalp area while maintaining local geometric consistency. The visualization is shown in Fig. 11.
Figure 11: Comparison w/ and w/o the post-processing procedure that enables smoother transition between the face and scalp parts modeled in unconnected regions of the UV space.
Appendix B Results
@@
Unconditional sampling. In Fig. 15, we provide more unconditional samples from our model from different viewpoints. All the samples have been produced by sampling $z\in\mathcal{N}(0,\mathbb{I})$ with a truncation trick [7] with the power $\psi=0.7$. For the evaluation in the Table 1 in the main text, the implementation of the metrics MMD, JSD, COV from PointFlow [59] was used. Since FaceVerse contains samples grouped by subjects, the nearest neighbor of a ground truth scan is typically a scan of the same subject with a different expression. Because of that, we only select one ground truth sample per subject (with the same neutral expression for all subjects) to calculate COV. All FaceVerse scans are used to calculate MMD and JSD. As a distance measure between individual point clouds, aggregated over multiple observations in MMD and COV, we use Chamfer Distance.
Ours Ours w/ VAE Ours w/ VQ-VAE
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 12: Ablation over the model design. VAE and VQ-VAE both follow the ResNet-18 encoder and decoder architecture, while our method is based on StyleGAN2. [VQ-]VAE mostly match the diversity of the training data but not the level of detail.
@@
Visual comparison of the generative model choices is demonstrated in Fig. 12.
Registration $L_{1}$-Chamfer $\downarrow$ N.C. $\uparrow$ F-Score @ 1.5 $\uparrow$
FLAME 3.33e-1 0.763 0.266
Stage 1 only 9.20e-2 0.817 0.861
Stage 1+2 (Ours) 5.68e-2 0.841 0.949
Table 3: Metrics reflecting the registration quality for our method w.r.t. only using Stage 1 of the registration procedure (vector offsets) and FLAME fits.
Behavior of the registration procedure. In Fig. 16, 17, 18, 19, we show how the mesh deforms as a result of the vector displacements optimization and normal displacements optimization.
Registration quality. In Table 3, we report exactly the same metrics as in NPHM [24] for both stages of our registration, averaged over identities. N.C. refers to Normal Consistency score; see [24] for clarification on all metrics. We excluded the region below the threshold by vertical axis (-0.30 in standard NPHM coordinate system) in order to not account for non-modeled region (clothes and neck). The registration precision is naturally better by all metrics than FLAME that does not provide a hair fit and better face fit. The necessity of the second stage is also motivated accordingly. Note that even though the numbers are not directly comparable to the ones in the NPHM paper [24], since there the evaluation over the face region only was conducted, and for us it is for the whole FLAME surface, they are still in the same scale. This indicates that the quality attained by non-rigid face refinement procedure in NPHM has been mostly achieved by our registration procedure for both face and hair.
@@
Comparison of the hair length. In Fig. 13, we show the comparison of the offset length in the scalp region between the ground truth data (NPHM scans) and the model predictions (HeadCraft). Since this offset length approximates the haircut size, we use it as a measure of hair length (not accounting for any hair curvature). This is demonstrated for comparison of the occurrence frequency of long hair samples in the model’s predictions w.r.t. the ground truth data distribution.
Figure 13: Comparison of the average offset length distribution in the scalp region for our method’s predictions and 3D scans from the NPHM dataset. We consider the offset length in the scalp region the approximation of the haircut size (or, simplified, a measure of hair length, without accounting for the hair curvature). The diversity of the average/maximum of this parameter per scan for our generated samples is on par with the training data. The horizontal axis quantifies the length in the NPHM coordinate system and the vertical axis stands for the bin height of the histogram.